### v1

@@ 2.4. Full irreducibility criterion
Suppose that $g\colon\Gamma\to\Gamma$ is a PNP-free, irreducible train track representative of $\varphi\in\textup{Out}(F_{r})$ such that $M(g)$ is Perron-Frobenius and all the local Whitehead graphs are connected. Then $\varphi$ is fully irreducible.
2.5. Principal fully irreducible outer automorphisms
@@
We include now a short description of properties of principal fully irreducible elements omitting definitions not used elsewhere in this paper. The interested reader can find the definitions in [MP16], which also includes a nice explanation of the history at the start of §2.9. For a principal $\varphi\in\textup{Out}(F_{r})$, the rotationless index is $i(\varphi)=\frac{3}{2}-r$. This implies $\varphi$ is ageometric and, by [MP16, Theorem 4.5], all its tt representatives are stable. In particular, none of its tt representatives has a PNP. By [MP16, Theorem 4.7] one has that $\varphi$ will also have only a single axis.
@@ 2.6. Folds & Stallings fold decompositions
[Uncaptioned image]Since they appear often, we describe here the notational conventions we use for proper full folds. Suppose that $f\colon\Gamma\to\Gamma^{\prime}$ is a single proper full fold of an edge $e_{1}$ over an edge $e_{0}$, as depicted to the right. Apart from $e_{1}$, each edge $e_{k}\in E\Gamma$ is mapped to a single edge of $\Gamma^{\prime}$, which we call $e_{k}^{\prime}$. The image of $e_{1}$ is an edge-path in $\Gamma^{\prime}$ consisting of 2 edges, the latter of which we call $e_{1}^{\prime}$. The map $f$ is then defined by
@@
and we just write $f\colon e_{1}\mapsto e_{0}e_{1}$. We call the edge-labeling of $\Gamma^{\prime}$ just described the induced edge-labeling. Note that it can be more convenient, especially where an orientation was chosen in advance, to write $f\colon e_{1}\mapsto e_{1}e_{0}$, where this was induced by folding $\bar{e}_{1}$ over $\bar{e}_{0}$. We may drop primes in complicated diagrams.
@@
In general, each fully irreducible $\varphi\in\textup{Out}(F_{r})$ has many train track representatives, each of which can have several distinct Stallings fold decompositions. By the main theorem of [MP16, Theorem 4.7], principal outer automorphisms (and their powers) each have a unique Stallings fold decomposition. Further, none of their train track representatives has a PNP.
@@ 2.7. The Outer space $\textup{CV}_{r}$ its (principal) geodesics
Given a Stallings fold decomposition of a train track map $g$, one can define a “periodic fold line” in $\textup{CV}_{r}$ using Skora’s [Sko89] interpretation of the decomposition as a sequence of folds performed continuously. In [AKKP19, Lemma 2.27] it is proved that periodic fold lines associated to train track maps are geodesics in the sense that, given 3 points $\gamma(t_{1}),\gamma(t_{2}),\gamma(t_{3})$ on the geodesic $\gamma$ with $t_{1}<t_{2}<t_{3}$, we have $d_{\textup{CV}_{r}}(\gamma(t_{1}),\gamma(t_{2}))+d_{\textup{CV}_{r}}(\gamma(t_{2}),\gamma(t_{3}))=d_{\textup{CV}_{r}}(\gamma(t_{1}),\gamma(t_{3}))$.
Since a general fully irreducible $\phi\in\textup{Out}(F_{r})$ can have several distinct Stallings fold decompositions, it can have several distinct periodic fold lines. However, the principal outer automorphisms that we are interested in will have just a single periodic fold line, its “lone axis.”
@@ 2.8. Fold-conjugate decompositions
Since an axis for a fully irreducible $\varphi\in\textup{Out}(F_{r})$ has a periodic structure, it becomes useful to view its Stallings fold decompositions cyclically. With some work one can see that starting at a different fold in a decomposition now yields a tt map representing an outer automorphism $\textup{Out}(F_{r})$-conjugate to $\varphi$ and with the same axis. It is also possible to start the tt map “in the middle of a fold.” We formalize here these different notions of cyclically permuting a Stallings fold decomposition or, equivalently, shifting along an axis.
[Uncaptioned image]A subdivided fold will mean a fold written as a composition of 2 folds, as depicted to the right. Suppose $\Gamma_{0}\xrightarrow{g_{1}}\Gamma_{1}\xrightarrow{g_{2}}\cdots\xrightarrow{g_{k-1}}\Gamma_{k-1}\xrightarrow{g_{k}}\Gamma_{k}$ and $\Gamma^{\prime}_{0}\xrightarrow{h_{1}}\Gamma^{\prime}_{1}\xrightarrow{h_{2}}\cdots\xrightarrow{h_{n-1}}\Gamma^{\prime}_{n-1}\xrightarrow{h_{n}}\Gamma^{\prime}_{n}$ are Stallings fold decompositions of homotopy equivalence tt maps $g\colon\Gamma\to\Gamma$ and $h\colon\Gamma^{\prime}\to\Gamma^{\prime}$. We say the decompositions are fold-conjugate if, possibly after a fold subdivision of $g_{k}$ into $\Gamma_{k-1}\xrightarrow{g_{k}^{\prime}}\Gamma_{k^{\prime}}\xrightarrow{g_{k+1}}\Gamma_{k}$ or $h_{n}$ into $\Gamma^{\prime}_{n-1}\xrightarrow{h_{n}^{\prime}}\Gamma^{\prime}_{n^{\prime}}\xrightarrow{h_{n+1}}\Gamma^{\prime}_{n}$, we have for some $j$ that one the following diagrams (1)-(3) commutes, where the vertical arrows are label-preserving graph isomorphisms.
@@
or
@@
or
@@
Fold-conjugate decompositions are partial-fold conjugate if we are in the case of (2) or (3) with $j=0$.
@@
We can also consider an equivalence relation generated by (a)-(c) on a more general class of fold sequences. This class of fold sequences would include pathological examples with many “unnecessary” subdivisions of one or more folds, as well as allowing us to start in the middle of any fold in the sequence. The benefit of such a relation is that it would allow one to define and understand an $\textup{Out}(F_{r})$-conjugacy class invariant that is the minimal number of folds in an equivalence class with respect to the partial-fold conjugacy relation.
@@ 3. The Rank-3 Principal Stratum Automaton $\widehat{\mathcal{A}_{3}}$
The “Lonely Direction Automaton” of [GP23] includes as directed loops only the Stallings fold decompositions of rotationless principal fully irreducible $\varphi\in\textup{Out}(F_{3})$. To include directed loops for all principal fully irreducible outer automorphisms, instead of just their rotationless powers, we construct a new directed graph we call the “Rank-3 Principal Stratum Automaton” $\widehat{\mathcal{A}_{3}}$. We construct $\widehat{\mathcal{A}_{3}}$ by adding in “permutation arrows” where ltt structure permutation relabelings exist in the following sense:
@@
Suppose $h\colon\Gamma^{\prime}\to\Gamma^{\prime}$ is a tt representative of $\varphi$ and let $R$ denote the rotationless power of $\varphi$. Then $h^{R}\colon\Gamma^{\prime}\to\Gamma^{\prime}$ is a tt representative of $\varphi^{R}$, and each $h$-periodic vertex and direction is fixed by $h^{R}$. By the proof of [MP16, Corollary 3.8], $h$ has at most 1 nonperiodic vertex and that vertex has precisely 2 gates. As in the proof of [MP16, Corollary 3.8], by performing a fold at that vertex, one obtains a tt representative $g^{\prime}$ of $\varphi^{R}$ for which all vertices are fixed and have at least $3$ fixed directions. Since $\varphi$ and $\varphi^{R}$ are lone axis fully irreducibles, this fold is within their shared axis $\mathcal{A}$.
@@
Because $\mathcal{A}$ is both $h$ and $h^{R}$-periodic, with both periods starting and ending at (re-marked copies of) $\Gamma^{\prime}$, the [MP16, Corollary 3.8] fold, let us call it $f$, is a fold of $\Gamma^{\prime}$ occurring at the start of the Stallings fold decompositions of both $h$ and $h^{R}$. If $f$ is a partial fold, since it occurs within the axis, there is some full fold $F$ in $\mathcal{A}$ and partial fold $f^{\prime}$ in $\mathcal{A}$ so that $F=f^{\prime}\circ f$. For similar reasons, $f^{\prime}$ would appear at the end of the Stallings fold decompositions. Hence, in fact, letting $g$ denote the conjugation of $h$ by the fold, $g^{\prime}=g^{R}$.
@@
We now show $g$ is a tt representative of $\varphi$. Since $g^{\prime}=g^{R}$ is a tt map, no power of $g$ can cause backtracking on an element of $E\Gamma$. Since quotienting by a partial fold cannot change the induced map of fundamental groups, we are left to show $g$ maps vertices to vertices. But each vertex has 3 $g^{\prime}$-gates, so cannot be mapped within its Stallings fold decomposition to a nonvertex point.
@@
Since $\varphi\in\textup{Out}(F_{r})$ is principal $\operatorname{IW}(\varphi)$, hence also $\operatorname{IW}(\varphi^{k})$ for each $k\in\mathbb{Z}_{>0}$, is the disjoint union of $32r-$ 3 triangles. By [MP16, Lemma 4.5] and since $\operatorname{IW}(\varphi)=\operatorname{IW}(\varphi^{k})$ has no cut vertices for each $k\in\mathbb{Z}_{>0}$, no tt representative $\tau$ of any $\varphi^{k}$ has a PNP. Thus, since $IW(\varphi^{k})$ is the disjoint union of the $SW(\tau,v)$ having at least 3 vertices, for any such $\tau$ and vertex $v$ with at least $3$ periodic directions, $SW(f,v)$ is a triangle. Since each vertex of $g^{R}$ has at least $3$ fixed directions, each vertex of $g$ has at least $3$ periodic directions. Hence, each $SW(g,v)$ is a triangle and there are $32r-$ 3 such triangles, proving (a).
@@
We now prove (b)-(c). Since each $SW(g,v)$ is a triangle, there are precisely 3 periodic directions at each vertex $v\in V\Gamma$. By [MP16, Lemma 3.6] $g^{R}$, hence $g$, has precisely one illegal turn. Thus, $\Gamma$ has precisely one $g$-nonperiodic direction. And we have that $\Gamma$ has a unique vertex of valence $4$, and all other vertices have valence $3$.
@@
(d) follows from [GP23, Lemma 2], after observing that each periodic direction must be contained in 2 turns (by the structure of the stable Whitehead graphs) so that the turn described in [GP23, Lemma 2] must contain the unique nonperiodic direction. ∎
@@
The Rank-3 Principal Stratum Automaton $\widehat{\mathcal{A}_{3}}$ is obtained from the “Lonely Direction Automaton” of [GP23] by adding a bi-directed edge labeled with $\sigma$ for each ltt structure permutation re-labeling $G\to\sigma\cdot G$, which implicitly records the directed edge $\sigma$ labeling the edge $[G,\sigma\cdot G]$ and $\sigma^{-1}$ labeling the edge $[\sigma\cdot G,G]$. For each $r\geq 3$, the Rank-r Principal Stratum Automaton $\widehat{\mathcal{A}_{r}}$ is defined analogously.
Figure 3 depicts $\widehat{\mathcal{A}_{3}}$ with the permutation edges depicted in green. While we refer the reader to [GP23] for the details of its construction, note that it is the maximal strongly connected components (in rank 3 there is only one such component containing a loop, up to permutation relabeling) of a graph where
  • vertices are ltt structures of the appropriate rank satisfying the “Lonely Direction Property” dictated by Lemma 3.2 and
  • directed edges are either folds compatible with the ltt structures at the endpoints, or permutation relabelings as discussed in our definition.
Following a directed loop thus gives a sequence of folds and permutations which could be a decomposition of a principal fully irreducible outer automorphism, though one can easily produce directed loops that do not give such elements.

### v2

@@ 2.4. Full irreducibility criterion
Suppose that $g\colon\Gamma\to\Gamma$ is a PNP-free, irreducible train track representative of $\varphi\in\textup{Out}(F_{r})$ such that $M(g)$ is Perron–Frobenius and all the local Whitehead graphs are connected. Then $\varphi$ is fully irreducible.
2.5. Principal & ageometric fully irreducible outer automorphisms
@@
We include now a short description of properties of principal fully irreducible elements omitting definitions not used elsewhere in this paper. The interested reader can find the definitions in [MP16], which also includes a nice explanation of the history at the start of §2.9. For a principal $\varphi\in\textup{Out}(F_{r})$, the rotationless index is $i(\varphi)=\frac{3}{2}-r$. This implies $\varphi$ is ageometric and, by [MP16, Theorem 4.5], all its tt representatives are stable. In particular, none of its tt representatives has a PNP ([BF94, Theorem 3.2]). By [MP16, Theorem 4.7] one also has that $\varphi$ has only a single axis (in the sense of §2.7).
@@ 2.6. Folds & Stallings fold decompositions
Since they appear often, we describe here the notational conventions we use for proper full folds. Suppose that $f\colon\Gamma\to\Gamma^{\prime}$ is a single proper full fold of an edge $e_{1}$ over an edge $e_{0}$, as depicted above. Apart from $e_{1}$, each edge $e_{k}\in E\Gamma$ is mapped to a single edge of $\Gamma^{\prime}$, which we call $e_{k}^{\prime}$. The image of $e_{1}$ is an edge-path in $\Gamma^{\prime}$ consisting of 2 edges, the latter of which we call $e_{1}^{\prime}$. The map $f$ is then defined by
@@
and we just write $f\colon e_{1}\mapsto e_{0}e_{1}$. We call the edge-labeling of $\Gamma^{\prime}$ just described the induced edge-labeling. Note that it can be more convenient, especially where an orientation was chosen in advance, to write $f\colon e_{1}\mapsto e_{1}e_{0}$ for the map induced by folding $\bar{e}_{1}$ over $\bar{e}_{0}$. We may drop primes in complicated diagrams.
@@
In general, each fully irreducible $\varphi\in\textup{Out}(F_{r})$ has many tt representatives, each of which can have several distinct Stallings fold decompositions. By [MP16, Theorem 4.7], principal outer automorphisms (and their powers) each have a unique Stallings fold decomposition.
@@ 2.7. The Outer space $\textup{CV}_{r}$ its (principal) geodesics
Given a Stallings fold decomposition of a tt map $g$, one can define a “periodic fold line” in $\textup{CV}_{r}$ using Skora’s [Sko89] interpretation of the decomposition as a sequence of folds performed continuously. In [AKKP19, Lemma 2.27] it is proved that periodic fold lines associated to tt maps are geodesics in the sense that, given 3 points $\gamma(t_{1}),\gamma(t_{2}),\gamma(t_{3})$ on the geodesic $\gamma$ with $t_{1}<t_{2}<t_{3}$, we have $d_{\textup{CV}_{r}}(\gamma(t_{1}),\gamma(t_{2}))+d_{\textup{CV}_{r}}(\gamma(t_{2}),\gamma(t_{3}))=d_{\textup{CV}_{r}}(\gamma(t_{1}),\gamma(t_{3}))$.
Since a general fully irreducible $\varphi\in\textup{Out}(F_{r})$ can have several distinct Stallings fold decompositions, it can have several distinct periodic fold lines. However, the principal outer automorphisms that we are interested in will have just a single periodic fold line, its “lone axis.”
@@ 2.8. Fold-conjugate decompositions
Since an axis for a fully irreducible $\varphi\in\textup{Out}(F_{r})$ has a periodic structure, it becomes useful to view its Stallings fold decompositions cyclically. With some work one can see that starting at a different fold in a decomposition now yields a tt map representing an element of $\textup{Out}(F_{r})$ that is $\textup{Out}(F_{r})$-conjugate to $\varphi$ and with the same axis. It is also possible to start the tt map “in the middle of a fold.” We formalize here these different notions of cyclically permuting a Stallings fold decomposition or, equivalently, shifting along an axis.
[Uncaptioned image]A subdivided fold will mean a fold written as a composition of a partial fold and then a completion of that fold, as depicted to the right. Suppose $\Gamma_{0}\xrightarrow{g_{1}}\Gamma_{1}\xrightarrow{g_{2}}\cdots\xrightarrow{g_{m-1}}\Gamma_{m-1}\xrightarrow{g_{m}}\Gamma_{m}$ and $\Gamma^{\prime}_{0}\xrightarrow{h_{1}}\Gamma^{\prime}_{1}\xrightarrow{h_{2}}\cdots\xrightarrow{h_{n-1}}\Gamma^{\prime}_{n-1}\xrightarrow{h_{n}}\Gamma^{\prime}_{n}$ are Stallings fold decompositions of homotopy equivalence tt maps $g\colon\Gamma\to\Gamma$ and $h\colon\Gamma^{\prime}\to\Gamma^{\prime}$, respectively. We say the decompositions are fold-conjugate if one of the following three possibilities occurs. In all cases the the vertical arrows are label-preserving graph isomorphisms.
Either the folds of the 2nd sequence are a cyclic shift of those of the 1st: there is some $j$ so that diagram (1) commutes.
@@
Or, the folds of the 1st are a cyclic shift of those of the 2nd, but starting midfold: there is some $\ell$ so that after subdividing $h_{\ell}$ into $\Gamma^{\prime}_{\ell-1}\xrightarrow{h_{\ell}^{\prime}}\Gamma^{\prime}_{{\ell}^{\prime}}\xrightarrow{h^{\prime\prime}_{\ell}}\Gamma^{\prime}_{\ell}$ diagram (2) commutes.
@@
Or, the folds of the 2nd are a cyclic shift of those of the 1st, but starting midfold: there is some $k$ so that after a fold subdivision of some $g_{k}$ into $\Gamma_{k-1}\xrightarrow{g_{k}^{\prime}}\Gamma_{k^{\prime}}\xrightarrow{g^{\prime\prime}_{k}}\Gamma_{k}$ diagram (3) commutes.
@@
Fold-conjugate decompositions are partial-fold conjugate if they fall in the case of (2) or (3), that is if one of the fold sequences is a cyclic shift of the other but for starting midfold. They are called partial-fold conjugate because the conjugating sequence contains a partial fold. In this case, $\mid n-m\mid=1$.
@@
We can also consider an equivalence relation generated by (1)-(3) on a more general class of fold sequences. This class of fold sequences would include pathological examples with many “unnecessary” subdivisions of one or more folds, as well as allowing us to start in the middle of any fold in the sequence. The benefit of such a relation is that it would allow one to define and understand an $\textup{Out}(F_{r})$-conjugacy class invariant that is the minimal number of folds in an equivalence class with respect to the partial-fold conjugacy relation.
@@ 3. The Rank-3 Principal Stratum Automaton $\widehat{\mathcal{A}_{3}}$
The “Lonely Direction Automaton” of [GP23] includes as directed loops only the Stallings fold decompositions of rotationless principal fully irreducible $\varphi\in\textup{Out}(F_{3})$. To include directed loops for all principal fully irreducible outer automorphisms, instead of just their rotationless powers, we construct a new directed graph we call the “Rank-3 Principal Stratum Automaton” $\widehat{\mathcal{A}_{3}}$. Before introducing $\widehat{\mathcal{A}_{3}}$, we review $\mathcal{A}_{3}$ as defined and constructed in [GP23].
3.1. The Lonely Direction Automaton $\mathcal{A}_{3}$
In [GP23], $\mathcal{A}_{3}$ is defined in a constructive manner using “comprehensive loops” and “permissible folds.” We give a simplified description of the construction now. $\mathcal{A}_{3}$ is the disjoint union of the strongly connected components of the following directed graph.
Nodes: There is a node for each order pair $(\Gamma,\ell)$ such that
$\bullet$ $\Gamma$ is a rank-3 graph (in the sense of §2) satisfying that all vertices are of valence-3, except a single vertex $v$ of valence-4 and
$\bullet \ell$ is a loop in $\Gamma$ traversing each edge of $\Gamma$ in at least one of the two directions and
$\bullet$ there is a direction $d$ at $v$ so that $\ell$ takes each turn of $\Gamma$ except precisely two of the turns at $v$ containing $d$.
The node is the ltt structure $ltt(\Gamma,\ell)$ defined in the same manner as in §2.3 except that the colored edges now come from the turns taken by $\ell$ and the single red vertex is that labeled by $d$. The nodes for $ltt(\Gamma,\ell)$ and $ltt(\Gamma^{\prime},\ell^{\prime})$ are equivalent (the same node), if there exists a decoration-preserving (meaning preserving both that color and labeling) graph isomorphism from $ltt(\Gamma,\ell)$ to $ltt(\Gamma^{\prime},\ell^{\prime})$. That is, the only recorded information from $\ell$ is the turns it takes.
Directed edges: There is a directed edge $E(f)$ from $ltt(\Gamma,\ell)$ to $ltt(\Gamma^{\prime},\ell^{\prime})$ precisely when there is a proper full fold $f\colon\Gamma\to\Gamma^{\prime}$ that does not fold any turn taken by $\ell$. That is, the turn folded by $f$ is not represented by a colored edge in $ltt(\Gamma,\ell)$.
3.2. The Rank-3 Principal Stratum Automaton $\widehat{\mathcal{A}_{3}}$
We build on $\mathcal{A}_{3}$ to construct $\widehat{\mathcal{A}_{3}}$ by adding in “permutation arrows” where ltt structure permutation relabelings exist in the following sense:
@@ 3.2. The Rank-3 Principal Stratum Automaton $\widehat{\mathcal{A}_{3}}$
Recall from §2.6 that for a proper full fold $f\colon\Gamma\to\Gamma^{\prime}$ we know that $|E\Gamma|=|E\Gamma^{\prime}|$ and $\Gamma^{\prime}$ comes with an induced edge-labeling. The induced edge-labeling then gives an identification of $E\Gamma$ with $E\Gamma^{\prime}$, and with this induced edge-labeling we can “push forward” the graph relabeling to $g_{\sigma}^{\prime}:\Gamma^{\prime}\to\sigma\cdot\Gamma^{\prime}$. Given a fold $f$ of $\Gamma$ and a graph relabeling $g_{\sigma}$, we obtain a fold of $\sigma\cdot\Gamma$ by “conjugating”: $f^{\prime}:=g_{\sigma}^{\prime}\circ f\circ g_{\sigma}^{-1}$. In what follows we will drop the prime from the notation $g_{\sigma}^{\prime}$, letting $g_{\sigma}$ refer to the permutation relabeling after pushing forward as many times as necessary in each context.
@@
Suppose $h\colon\Gamma^{\prime}\to\Gamma^{\prime}$ is a tt representative of $\varphi$ and let $R$ denote the rotationless power of $\varphi$. Then $h^{R}\colon\Gamma^{\prime}\to\Gamma^{\prime}$ is a tt representative of $\varphi^{R}$, and each $h$-periodic vertex and direction is fixed by $h^{R}$. By the proof of [MP16, Corollary 3.8], $h^{R}$, and hence also $h$, has at most 1 nonperiodic vertex and that vertex has precisely 2 gates. As in the proof of [MP16, Corollary 3.8], by performing a fold at that vertex, one obtains a tt representative $g^{\prime}$ of $\varphi^{R}$ for which all vertices are fixed and have at least $3$ fixed directions. Since $\varphi$ and $\varphi^{R}$ are lone axis fully irreducible outer automorphisms, this fold is within their shared axis $\mathcal{A}$.
@@
Because $\mathcal{A}$ is both $h$- and $h^{R}$-periodic, with both periods starting and ending at (re-marked copies of) $\Gamma^{\prime}$, the [MP16, Corollary 3.8] fold, let us call it $f$, is a fold of $\Gamma^{\prime}$ occurring at the start of the Stallings fold decompositions of both $h$ and $h^{R}$. If $f$ is part of a subdivided fold (in the sense of §2.8), since it occurs within the axis, there is some full fold $F$ in $\mathcal{A}$ and partial fold $f^{\prime}$ in $\mathcal{A}$ so that $F=f\circ f^{\prime}$. When you do the folds cyclically, since the decomposition of $h$ lies on $\mathcal{A}$, the copy of $f^{\prime}$ at the end of the decomposition of $h$ must compose with $f$ at the beginning of the next iteration of $h$ to give $F$. Thus, $f^{\prime}$ would appear at the end of the Stallings fold decompositions of $h$ and $h^{R}$. Hence, in fact, letting $g$ denote the conjugation of $h$ by the fold, i.e. $f\circ h=g\circ f$, we have $g^{\prime}=g^{R}$.
@@
We now show that $g$ is a tt representative of $\varphi$. Since $g^{\prime}=g^{R}$ is a tt map, no power of $g$ can cause backtracking on an element of $E\Gamma$. Since quotienting by a partial fold cannot change the induced map of fundamental groups, we are left to show that $g$ maps vertices to vertices. But each vertex has at least 3 $g^{\prime}$-gates, so cannot be mapped within its Stallings fold decomposition to a nonvertex point.
@@
Since $\varphi\in\textup{Out}(F_{r})$ is principal, its ideal Whitehead graph $\operatorname{IW}(\varphi)$, hence also $\operatorname{IW}(\varphi^{k})$ for each $k\in\mathbb{Z}_{>0}$, is the disjoint union of $32r-$ 3 triangles. By [MP16, Lemma 4.5] and since $\operatorname{IW}(\varphi)=\operatorname{IW}(\varphi^{k})$ has no cut vertices for each $k\in\mathbb{Z}_{>0}$, no tt representative $\tau$ of any $\varphi^{k}$ has a PNP. Thus, since $IW(\varphi^{k})$ is the disjoint union of the $SW(\tau,v)$ having at least 3 vertices, for any such $\tau$ and vertex $v$ with at least $3$ periodic directions, $SW(\tau,v)$ is a triangle. Since each vertex of $g^{R}$ has at least $3$ fixed directions, each vertex of $g$ has at least $3$ periodic directions. Hence, each $SW(g,v)$ is a triangle and there are $32r-$ 3 such triangles, proving (a).
@@
We now prove (b)-(c). Since each $SW(g,v)$ is a triangle, there are precisely 3 periodic directions (hence gates) at each vertex $v\in V\Gamma$. By [MP16, Lemma 3.6], $g^{R}$, hence also $g$, has precisely one illegal turn. Thus, $\Gamma$ has precisely one $g$-nonperiodic direction. And we have that $\Gamma$ has a unique vertex of valence $4$, and all other vertices have valence $3$.
@@
Part (d) follows from [GP23, Lemma 2], after observing that each periodic direction must be contained in 2 turns (by the structure of the stable Whitehead graphs) so that the turn described in [GP23, Lemma 2] must contain the unique nonperiodic direction. ∎
@@
The Rank-3 Principal Stratum Automaton $\widehat{\mathcal{A}_{3}}$ is obtained from the “Lonely Direction Automaton” of [GP23] by adding a bi-directed edge labeled with $\sigma$ for each ltt structure permutation relabeling $G\to\sigma\cdot G$, which implicitly records the directed edge $\sigma$, labeling the edge $[G,\sigma\cdot G]$, and $\sigma^{-1}$, labeling the edge $[\sigma\cdot G,G]$. For each $r\geq 3$, the Rank-r Principal Stratum Automaton $\widehat{\mathcal{A}_{r}}$ is defined analogously.
Note that, once a choice of labels has been made for the black edges of a single node, this determines the labeling of the black edges in each other node, as well as the labels on the directed edges. In order to explicitly write out $\widehat{\mathcal{A}_{3}}$, we make such a choice and call the decoration-preserving graph isomorphism class of $\widehat{\mathcal{A}_{3}}$, endowed with such a choice of labeling, a graph-labeling class.
Figure 3 depicts $\widehat{\mathcal{A}_{3}}$ with the permutation edges depicted in green. The nodes and black directed edges are as described in §3.1. (Strictly speaking, what is shown is a maximal strongly connected component of $\widehat{\mathcal{A}_{3}}$; in rank 3 there is only one such component containing a loop, up to permutation relabeling.) Following a directed loop thus gives a sequence of folds and permutations which could be a decomposition of a principal fully irreducible outer automorphism, though one can easily produce directed loops that do not give such elements. For example, the folds represented by the directed edges of the loop may be supported in a proper subgroup of the fundamental group (so that the map will be reducible).
3.3. Loops in $\widehat{\mathcal{A}_{3}}$